### v1

# SpiderGen:Towards Procedure Generation For Carbon Life Cycle Assessments with Generative AI

Abstract
Investigating the effects of climate change and global warming caused by GHG emissions have been a primary concern worldwide. These emissions are largely contributed to by the production, use and disposal of consumer products. Thus, it is important to build tools to estimate the environmental impact of consumer goods, an essential part of which is conducting Life Cycle Assessments (LCAs). LCAs specify and account for the appropriate processes involved with the production, use, and disposal of the products. We present SpiderGen, an LLM-based workflow which integrates the taxonomy and methodology of traditional LCA with the reasoning capabilities and world knowledge of LLMs to generate the procedural information used for LCA. We additionally evaluate the output of SpiderGen using real-world LCA documents as ground-truth. We find that SpiderGen provides accurate LCA process information that is either fully correct or has minor errors, achieving an F1-Score of 62% across 10 sample data points. We observe that the remaining missed processes and hallucinated errors occur primarily due to differences in detail between LCA documents, as well as differences in the “scope” of which auxiliary processes must also be included. We also demonstrate that SpiderGen performs better than several baselines techniques, such as chain-of-thought prompting and one-shot prompting. Finally, we highlight SpiderGen’s potential to reduce the human effort and costs for estimating carbon impact, as it is able to produce LCA process information for less than \$1 USD in under 10 minutes as compared to the status quo LCA, which can cost over \$25000 USD and take up to 21-person days.
1 Introduction
Figure 1: A simplified example PFG for the product category “Wine” produced by SpiderGen.
Figure 2: The SpiderGen workflow utilizes LLMs and sentence transformers to create Product Category Rule process flow graphs. This workflow consists of (1) A product generation step, where sample products are sourced and (2) processes for those sample products are generated, (3) A coarse process generation step, where the sample product processes are coarsened to create generalized processes and (4) the generation of the process flow graph itself. The formation of these graphs can then be used for a variety of applications, such as carbon footprint and supply chain analysis.
Life Cycle Assessments (LCAs) are commonly used to evaluate the environmental impact of a product in all phases of the product’s “life”, such as the manufacturing phase or the use phase by the end consumer. Although LCAs are essential for measuring the environmental impacts of products, the end-to-end time and monetary cost of conducting LCAs is extraordinarily high - a 2017 study shows that just determining the appropriate processes for a given product category can cost over to \$25000 USD, and can take up 21 person days to produce (tasaki_international_2017). Due to the challenges of generating the appropriate processes for an LCA, it is very difficult to scale environmental impact assessments. As a result, LCAs are only available for a very small subset of products.
A key challenge in creating an LCA is choosing which processes to include in the analysis, and creating standardized methodologies with which to evaluate specific categories of products. Traditionally, the appropriate processes for a given product category are determined by a committee of experts, and the LCA practitioner utilizes these processes and augments them for the specific product. These processes are described using a Process Flow Graph (PFG) to represent the ordering and dependencies throughout the value chain of the product. An example of a PFG for the product category “Wine” is shown in Figure 1.
To address these challenges, there has been a growing interest in using machine learning techniques to automate LCA procedures to allow for easier, faster and more ubiquitous LCAs. Prior work has primarily focused on using existing process rules for specific categories of products, “bill of materials” (BOMs), and LCA guidelines, and automating other pieces of the LCA pipeline (balaji_caml_2023; balaji_flamingo_2023; balaji_parakeet_2024; zhang_deltalca_2024; sousa_approximate_2002; Wang2025). However, to our knowledge, we are the first to explore the potential of using ML to generate PFGs for LCA. Automating this key step will enable LCAs of products which have no guiding PFG available previously, such as laptops. This problem is complex, as it requires the collection of a wide range of process information, understanding the dependencies of these processes, and the ability to generalize processes to be applicable to all products of a given category.
Our Approach. We propose SpiderGen  [^2] Much like a spider, SpiderGen is a workflow which generates a web of resources and weaves insights together to generate PFGs., a novel LLM-based workflow that automates the production of PFGs for a given product category. To do so, SpiderGen utilizes the world knowledge capabilities of LLMs, as well as text-embedding methods such as SBERT (reimers-2019-sentence-bert), and graph clustering algorithms such as K-Means clustering to derive upstream, core, and downstream processes for the overarching product category. SpiderGen then orders these processes to produce a Directed Acyclic Graph (DAG) representation of the end-to-end life cycle of the product category. Notably, we leverage specific ontologies and taxonomies involved in LCA production, to be able to generate relevant processes that are required in the PFG. We specifically address the challenges of generating processes that are generalizable across all products of a given product category, ordering them based on their dependencies, and ensuring that all of the processes that may be within the “scope” of a true LCA are included.
Evaluation Methodology. To evaluate SpiderGen, we utilize a ground truth dataset of 66 LCA documents from EPD International (EPD). To compare the generated PFGs to the ground truth, we introduce a novel evaluation methodology (Section 4) since prior knowledge graph evaluation methods, such as those for UML diagrams and Buisness Process Flow graphs are not appropriate for our problem (fauzan_structural_2024; faisal_approach_2018; kristina_business_2024). Complete PFGs must include all the correct processes and nothing more, and must include the correct partial ordering of processes, which reflect the dependencies of each process. To determine whether SpiderGen has generated the correct processes and ordering, we provide both a qualitative and quantitative evaluation of SpiderGen. For qualitative evaluation, we outline a set of criteria that determine “correctness” and evaluate 10 datapoints using this criteria. For the quantitative evaluation, we calculate the Pointwise Mutual Information (PMI) between the ground-truth and SpiderGen’s generated PFGs.
Our Contributions. SpiderGen serves as a key tool to determine the environmental impacts of many products. We make the following contributions:
  • We describe and implement SpiderGen, a novel zero-shot workflow for producing PFGs which utilizes the world-knowledge and text processing capability of LLMs.
  • We introduce an evaluation methodology for comparing PFGs using Pointwise-Mutual Information estimation, and conduct a qualitative analysis of SpiderGen’s ability to produce the necessary processes for LCA.
  • We implement and apply the SpiderGen pipeline to generate PFGs using the United Nations Central Product Classification (UNCPC) descriptions and evaluate the generated PFGs using real-world LCA procedure documents for general product categories.
  • We find that SpiderGen provides accurate PFGs as compared to the ground truth, producing processes that are fully correct or only with minor errors 62% of the time on average. We find that the main challenges of increasing the accuracy of SpiderGen lie in producing all of the appropriate auxiliary processes that are part of the ground truth PFG, while avoiding extrapolated processes that may not be applicable to a given product category.
2 Related Work
Traditional LCA Frameworks
LCA is the method used to estimate the carbon footprint of products and processes. A LCA expert starts by identifying the processes and environmental factors of the product life cycle that must be included. To do so, LCA experts rely on Process Flow Graphs (PFGs), which describe the processes for a specific product category that must be included in the LCA and their dependencies. By utilizing these graphs, experts can avoid making errors while conducting LCAs. Further, these graphs allow products of the same category to be evaluated in a consistent way, allowing for appropriate LCA comparisons. Additionally, these graphs also prevent “greenwashing”, where environmental impact information is manipulated to make products appear more sustainable than they truly are.
PFGs are officially made available within Product Category Rules documents, and are created by a team of industry experts. However, the process of creating PFGs is time and resource intensive, making them inaccessible. For example, although PFGs exist for many food products, they do not exist for computing technology products, such as laptops. Due to the lack of accessible LCA information, there is growing interest in automating LCA using ML.
Machine Learning For LCA
Incorporating machine learning techniques for different parts of the LCA process has become a growing area of interest (algren2021).
Some of the earlier efforts utilize surrogate modeling to create new LCA analyses based on a training set of LCAs for a set number of products (sousa_approximate_2002; sousa_learning_nodate; sousa_product_2006).
More recently, researchers have proposed using machine learning and LLMs to extract relevant information from LCA reports and databases (goridkov_whats_2024). Balaji et al. use sentence embeddings and LLMs to match environmental impact factors from LCA databases to a given product descriptions (balaji_flamingo_2023; balaji_caml_2023; balaji_parakeet_2024). A related body of work uses information extraction methods to synthesize LCAs by utilizing templates for a given product, such as a Bill of Materials (BOM) and domain-specific databases to guide the extraction, synthesis or prediction of LCAs (zhang_deltalca_2024). In contrast to our work, these prior works do not explore the generation of PFGs for LCA.
Procedural Generation using LLMs
Using LLMs to generate procedural information has also become popular in different domains such as generating food recipes, UML diagrams, and Business Process Flow diagrams. The approaches either generate new procedures using a fine-tuned language model (Mohbat_2024), or extract structured outputs from existing procedure documents. The most closely related work to ours utilizes LLMs to structure the outputs of existing Standard Operating Procedures in the case of recipe generation and Business Process Flow diagrams (garg2025generatingstructuredplanrepresentation). While this work organizes existing procedural information into a standard operating procedure graph, our approach contains an additional step of generating the processes themselves.
Our Approach
To the best of our knowledge, we are the first to explore the problem of generating process flow graphs for LCA. In contrast to prior work, our approach is zero-shot and uses off-the-shelf models to generate PFGs. We achieve this by applying the ontology and taxonomies of LCA production to create an LLM workflow for graph construction for our specific problem. Unlike fine-tuning based methods, SpiderGen is not constrained by limited training data, which allows SpiderGen to create a PFG for almost every product category.
3 Problem and Preliminaries
We introduce SpiderGen, an LLM-based workflow which generates a graphical representation of the life cycle processes of a given product, called a Process Flow Graph (PFG). We define the PFG for a given product category $pc$ to be $G_{pc}=(V_{pc},E_{pc})$, where the vertices $V_{pc}$ represent the generalized processes for a given product category, and the edges $E_{pc}$ represent their ordering. $G_{pc}$ generation uses the ontology and taxonomy of LCA PFG construction. The nodes in $V_{pc}$ span three life cycle phases: “upstream” processes include all life cycle processes that come before the manufacturing of the product (ex. raw material extraction), “core” processes involve all manufacturing steps, and “downstream” processes include processes that come after product manufacturing (ex. consumption). To reflect this taxonomy, each node in $V_{pc}$ is given a label for the life cycle phase during which it occurs (either “upstream”, “core” or “downstream”). In our setting, $G_{pc}$ is a directed acyclic graph (DAG) and is made up of two sets of edges: $E_{m}$ and $E_{s}$, such that $(E_{pc}=E_{m}\cup E_{s}$). $E_{m}$ indicates the ordering between main processes (i.e, for edge $(i,j)\in E_{m}$, process $i$ precedes process $j$). $E_{s}$ indicates the relationship between sub-processes and main processes (i.e., for edge $(i,j)\in E_{s}$, $i$ is a subprocess of j). Additionally, we require that all upstream processes must precede core processes, and all core processes must precede downstream processes.
4 SpiderGen: Our Proposed Solution
We now describe how SpiderGen produces $G_{pc}$. Generating $G_{pc}$ involves the following steps (as shown in Figure 2).
(1) Generate Sample Products: Since $G_{pc}$ is generalizable across different products in a given category, we generate a list of real products that are relevant to this category. The goal of this step is to get a lay of the land of common processes required for products in this category, drawing this information as much as possible from real examples. There are two challenges with generating sample products:
Ensuring Product Diversity: The listed products must ideally use a diverse set of processes and raw materials to avoid the issue of overfitting to an overrepresented type of product within a product category.
Quantity: To gain a sufficient understanding of a product category, we must ask for the right number of products to use. In early experiments, we found that asking for too few products would result in PFG information that was over-fitted to a specific niche within a product category. However, querying for too many products may make the PFG over-generalized and therefore inaccurate.
(2) Generate Sample Product Process Lists: To find common processes across a given product category, SpiderGen instructs the LLM to select diverse products that use a wide variety of materials and processes. For each product, SpiderGen queries the LLM for details about what components make up the product, and how these components are processed. We then query the LLM to find the processes involved in manufacturing and distributing the product. We prompt the LLM to list the process name, the life cycle phase it is involved in, and describe why this process is included. We use the LCA ISO standards to guide the LLM for process generation to cover various types of processes and ensure complete coverage of environmental impacts (EPD).
(3) Generate Coarse Process List: From the detailed process descriptions of a list of products, we create coarser generalized processes that can be included in a PFG. Intuitively, similar processes that appear in all product processes should be included in the PFGs, since it is likely that they are an important process to include for the entire category. Based on this intuition, we identify clusters of common processes to include in the final PFG. Finally, we utilize an LLM to summarize these clusters to create coarse process descriptions and eliminate clusters that are repeated. To generate relevant clusters of nodes, we utilize a pre-trained SBERT model to embed each process. We then use K-means clustering to provide groupings of repeated processes. To ensure that we do not cluster processes that may seem similar but are a part of different life cycle phases within the final PFG, we only cluster processes that are in the same life cycle phase. For each life cycle phase, we select the number of clusters that minimizes the Davies Bouldin score of each group life cycle phase clusters. Using these clusters, we then prompt the LLM to give a description of each of the clusters and then remove repeated descriptions. The resultant descriptions are the coarse processes used in the final process flow graph.
(4) Generate the final PFG: Given the processes that form the vertices of $G_{pc}$, we now generate the edges of the graph to place these processes in the correct ordering. We establish two types of ordering: explicit ordering and implicit ordering. Implicit ordering is based on the life cycle phases associated with each process. For example, processes that are a part of the “upstream” phase will always appear before processes that are a part of the “core” phase. After creating these implicit orderings, we use an LLM to generate explicit orderings within processes with the same life cycle phase. For example, the LLM will reason that a process called “Mixing dough” will appear before “Forming noodles with dough”, where both are core processes. From our observation, modern LLMs can reason about explicit ordering quite well, and provide orderings that align with real process orderings. By using implicit ordering based on life cycle phases dictated by the taxonomy of LCA, SpiderGen is able to generate more accurate explicit orderings within a smaller set of processes that belong to the same life cycle phase. Using this method, we observe that SpiderGen is able to create process orders that are close to the correct PFGs.
5 Evaluation Methodology
Evaluation Setup
We evaluate the generation of PFG $G_{pc}$ on a set of 66 Product Category Rules (PCR) documents in the EPD International Database (EPD). These documents contain generalized processes for a given product category or a set of product categories for each phase of the product life cycle (upstream, core, and downstream). We manually extract the processes from these PCR documents and form a PFG with these processes. [^3] We initially attempted to use Retrieval Augmented Generation (RAG) to extract process descriptions from the PCR documents. However, due to the lack of consistency across different documents, the RAG sometimes missed smaller processes. To ensure full correctness in the ground truth, we opted to conduct this extraction manually. We provide the code for the RAG method in our GitHub repository. Each of the PCR documents have United Nations Central Product Classification (UN CPC) codes that describe the product category. For each PCR, we compile descriptions of the relevant UN CPC codes, as well as the name of the PCR. This UN CPC information is given to SpiderGen as an input for generating the PFG. We experiment with three different LLM models: OpenAI’s GPT-4o, o1-preview and o1-mini models. We used these models for our evaluation in April, May, June and July of 2025.
Baseline Methods for Evaluation
As there are no existing evaluation baselines for our problem, we present two baseline methods which draw upon existing techniques in LLM prompt engineering:
LLMDirect: The LLMDirect method directly prompts the LLM for the PFG using the well-established method of Chain-of-Thought Reasoning (sahoo2025systematicsurveypromptengineering). With this method, we provide the LLM with step-by-step instructions on how to gather all the necessary processes to create the PFG, as well as how to create the graph using these steps. We also prompt it to provide a rationale for each step.
LLMExample: The LLMExample method is an improvement on LLMDirect and involves the developer conducting One-Shot Prompting (sahoo2025systematicsurveypromptengineering), where the LLM is given an example PFG, and must follow this example to create a new PFG for an entirely different product category.
Qualitative Evaluation Methodology
We qualitatively analyze the results for 10 different product categories by comparing the PFGs generated by SpiderGen to the ground truth PFGs, categorizing any errors that SpiderGen makes. This procedure involves two steps: We first pair all of the nodes by similarity by pairing every node from the generated graph $G_{pc}$ to the most similar node from the ground truth graph. Every node from both graphs must be included in the matching, and many-to-many matching are allowed. If there is no appropriate pairing for a given node, we pair the node with a fake node called ”N/A”.
We then evaluate the matches using the following criteria:
  • If there exists a match such that the ground truth node corresponds to that in the generated graph, we label this pairing as a “match”. Note that we are considering cases where part of the process label is exactly matching to be labeled as “match” (i.e, if there is a ground-truth node called “Agriculture, including electricity and water consumption” and there is a generated node called “electricity consumption of agricultural processes”, this is a “match”).
  • If there exists a generated node that is an inferred (but not explicitly stated) sub-process of a ground-truth node, this match is labeled as “subprocess” (i.e, if there is a ground-truth node called “Manufacturing Pasta” and a generated node called “Cutting pasta dough”, this is a “subprocess”)
  • If there exists a generated node that is a specific version of a process stated in the ground-truth graph, this match is labeled as “specific”. (i.e, if there is a ground-truth node called “Cultivating grains for pasta”, and there is a generated node called “cultivating durum wheat for pasta”, this is a “specific process”).
  • If there exists a generated node which does not have an appropriate pairing with a ground-truth graph node, it is labeled as “wrong”, because in these scenarios, SpiderGen has hallucinated a step that does not exist for the product category. If there exists a ground-truth node which does not have an appropriate generated node pair, it is labeled as “missing”, because in these scenarios, SpiderGen has missed including the process.
F1 Scoring: We calculate an F1 score to capture the overall quality of the generated graph. The F1 Score is defined as
$(2*\text{Precision}*\text{Recall})/(\text{Precision}+\text{Recall})$ (1)
where Recall is calculated as
$\text{Recall}=(c_{specific}+c_{subprocess}+c_{match})/(n_{groundtruth})$ (2)
and Precision is calculated as
$\text{Precision}=1-(c_{wrong})/(n_{generated})$ (3)
where $c_{specific}$ is the count of ground truth nodes that are in “specific” pairings, $c_{subprocess}$ is the count of ground-truth nodes that are in “subprocess” pairings, $c_{match}$ is the count of ground truth nodes that are in “match” pairings, $n_{groundtruth}$ is the total number of ground truth nodes, $c_{wrong}$ is the count of generated nodes that are labeled as “wrong”, and $n_{generated}$ is the total number of generated nodes.
We provide a small sample of manually evaluated [^4] We initially utilized LLM-as-a-Judge techniques to rapidly conduct multiple of these evaluations. However, we observed that even with the addition of example cases, and chain-of-thought prompting, LLM-as-a-Judge tended to estimate a higher recall than a human judge. analyses, and describe themes in the erroneous processes produced by SpiderGen. This qualitative analysis for 10 product categories is given in Table 1. We observe that the majority of errors produced by SpiderGen derive from two sources: (1) misalignment of an appropriate “scope” for the processes that must be included (i.e, how many tangential or auxiliary processes should be included), and (2) extrapolation and assumptions about how a product might be used and what materials may be involved to create the product.
Quantitative Evaluation Methodology
To evaluate node similarity between the generated and the ground truth graphs, we calculate the Pointwise Mutual Information (PMI) to determine the similarity between lists of processes (xu_benchmarking_2025). Pointwise mutual information is a commonly used measure to compare the similarity of two texts, and measures the probability of two texts co-occuring by chance. In the case of text-comparison tasks, PMI is defined as
$PMI(X=x,Y=y)=\log{\Pr(Y=y|X=x)}-\log{\Pr(Y=y)}$ (4)
where $X$ and $Y$ are two texts.
In our case, we calculate the PMI across two lists of processes, one from the SpiderGen generated $G_{pc}$, and the other from a ground truth LCA document . We consider $X$ to represent the random variable for the ordered list of processes generated by SpiderGen, and $Y$ to be the random variable representing the ordered list of ground truth processes.
Prior work indicates that using LLM-weights to calculate PMI can lead to improved evaluations, as this metric can be more sensitive to semantic degradations when comparing the semantic similarity of two texts (xu_benchmarking_2025). Thus, we closely follow this methodology from prior work. We utilize the Llama-8b-Instruct model weights for all texts for consistency, and calculate a normalized PMI using the LLM weights as our quantitative evaluation metric.
For our quantitative analysis, we follow the steps below:
  • Pre-processing both lists: We organize both lists such that upstream processes are listed first, core processes second, and downstream processes third. Both lists start with “raw material procurement” processes, and end with “end-of-life” processes. For ground-truth documents, direct references to other documents, links, references to other sections of the text and references to calculations have been removed. Additionally, processes marked as “optional” are marked as such. Xu et.al propose “pre-processing” text by using an LLM to rephrase the texts before calculating PMI. However, this pre-processing is not feasible for our problem as it may lead to a higher risk of introducing hallucinations that can change the meaning and structure of the ground truth text.
  • Generating Log Probabilities: To generate the log probabilities for $\log{\Pr(Y=y|X=x)}$, we create a combined text block, first listing the ground-truth node list and subsequently adding the generated node list. We then tokenize the entire text block and generate the probabilities for the sequence of text using the weights of Llama-8b-Instruct. We then calculate the log of these probabilities. Similarly, we calculate $\log{\Pr(Y=y)}$ by tokenizing the list of ground-truth nodes and getting the probabilities via the weights of Llama-8b-Instruct. By calculating both $\log{\Pr(Y=y|X=x)}$ and $\log{\Pr(Y=y)}$), we can calculate the PMI of any two lists.
  • Normalizing the PMI: To normalize the PMI, we calculate the maximum possible PMI of the list of nodes. The maximum possible PMI occurs when the exact same text is generated twice. Thus, we calculate $PMI(x,y)/PMI(y,y)$, given that $y$ is the ground-truth node list, and $x$ is the generated node list.
Using this method, we are able to provide a quantitative evaluation whether the content of the process lists are similar between ground-truth process lists and SpiderGen’s generated process lists.
Product Category PMI F1-Score
Railways (44) 0.02 0.37
Shower Enclosures (40) 0.02 0.66
T-Shirts, Tops (28) 0.03 0.63
Moka Coffee (22) 0.02 0.63
Dairy Products (22) 0.05 0.61
Graphite Products (20) 0.04 0.63
Detergents & Washing (19) 0.09 0.50
Woven Fabric (17) 0.05 0.84
Air Ducts (17) 0.07 0.71
Bottled Water (12) 0.03 0.59
Table 1: Results of our Qualitative Analysis for 10 products (Shown circled in Figure 3). We order the products based on complexity denoted by the number of nodes (in parenthesis). We observe higher F1 scores for products with higher PMI values in general. Note that PMI is the quantitative metric defined in Equation 4.
Figure 3: Comparing the normalized PMI values as different numbers of products in each category are generated in the first step of the SpiderGen workflow. We observe lower variability in PMI as the number increases.
Figure 4: Comparing the Normalized PMI Scores for 66 products with varying complexity as denoted by the number of nodes in the ground truth $G_{pc}$ graphs. SpiderGen has higher normalized PMI for simpler product categories (e.g. “Dairy Products”) and lower for more complicated ones (e.g. “Railways”). We also report the qualitative scores for 10 different product categories (circled nodes) in Table 4 .
6 Evaluation Results
We present an evaluation of SpiderGen, highlighting key results.
Model Average PMI Average Cost per PFG
o1-preview $0.051\pm 0.005$ US\$ $4.7\pm 0.28$
o1-mini $0.046\pm 0.006$ US\$ $0.36\pm 0.03$
gpt-4o $0.039\pm 0.003$ US\$ $0.64\pm 0.14$
Table 2: Average PMI Score and cost for SpiderGen using different LLMs. SpiderGen performs better with models, such as OpenAI’s o1-preview and o1-mini. o1-mini provides a good tradeoff with good performance and lowest cost.
Evaluating the Parameters of SpiderGen
We first examine two key parameters that affect the SpiderGen workflow: (a) the selection of LLM model and (b) the number of example products used in the initial product generation step.
Efficacy vs Cost trade-offs of utilizing reasoning models:
We implemented SpiderGen using 3 OpenAI models: gpt-4o, o1-mini, and o1-preview. o1-mini and o1-preview are reasoning models, with o1-mini being a smaller, faster, and cheaper model than the other two. Table 2 compares the performance of these models in terms of the average normalized PMI and cost for generating a single PFG. The two reasoning models (o1-mini and o1-preview) outperform gpt-4o since they are able to provide more detailed outputs. However, o1-mini is close in terms of performance to o1-preview while being 1/13th the cost at less than US \$1 on average per PFG. This cost is far lower than the typical production of a PFG, which can be upwards of US \$25000.
Effect of the number of products generated by SpiderGen:
We evaluate the effect of producing a different number of sample products in the first step of SpiderGen (shown in Figure 3 for four product categories). We evaluate the production of PFGs using 3, 6, 9, 12, 15 sample products for each product category. As shown in the figure, the PMI often increases with the addition of more sample products. We note that less “niche” product categories, such as “Grain Mills” benefit from a larger number of sample products, and more “niche” categories, such as “Moka Coffee” benefit from smaller numbers of sample products, and can even have deteriorated PFGs, as the number of sample products increases. Since generating 15 products works well for many scenarios, we use it for our implementation of SpiderGen.
Evaluating SpiderGen on Ground Truth Data
To evaluate SpiderGen end-to-end, we execute our pipeline for 66 product categories and compare the generated PFGs to the ground truth graphs extracted from real LCA documents (EPD).
Efficacy of SpiderGen compared to baselines:
We compared SpiderGen with the two baseline methods LLMDirect and LLMExample across 40 product categories. Our resuts show that SpiderGen exceeds both baseline methods, providing a median normalized PMI score of $0.043\pm 0.025$ across different product categories as compared to $0.026\pm 0.022$ for LLMDirect and $0.024\pm 0.024$ PMI for LLMExample. Based on our qualitative analysis of a subset of ten products, we find that SpiderGen avoids generating overly specific processes and captures more non-obvious auxiliary processes as compared to the baselines.
Limitations of LLMDirect: Although a Chain-of-Thought process may work for many problems, we found that an algorithm such as LLMDirect is insufficient to solve our problem. Firstly, it is unable to capture all the nuanced steps that are necessary for complete PFGs. LLMDirect results frequently misses steps that are peripheral to the main manufacturing processes. Further, it is unclear whether the resultant PFG is truly applicable to all products in a specific product category and is likely to be overly specific to one subcategory of the product category.
Limitations of LLMExample: While LLMExample is more effective than LLMDirect, since it is able to capture the nuanced steps of the PFG by following an example, it is still insufficient to solve our problem. LLMExample misses peripheral steps that may be important to a broader product category. Additionally, the PFG is still overly specific to a subset of items of the product category. This is particularly likely if the example product category is more niche than the product category information that we are generating.
SpiderGen on Varying Product Complexity:
PFGs can be described as more complex if they have more life cycle processes. For example, “railways” is the second-most complicated product category in our dataset, as the PCR for “railways” includes 44 different processes that must be taken into account. To evaluate whether SpiderGen is able to produce these complex PFGs, we evaluated SpiderGen on ground-truth PFGs with varying levels of complexity. Figure 4 shows a scatter plot of the normalized PMIs of PFGs with different numbers of nodes. We find an overall trend is that for products that require more nodes, the PMI is lower than with product categories with smaller PFGs. Both of these factors indicate that SpiderGen is sensitive to a change in product complexity. As it stands, SpiderGen produces higher-quality PFGs for simpler products.
7 Discussion and Conclusion
In this work, we introduce SpiderGen, a novel machine learning workflow that automates the generation of PFGs, and introduce evaluation methods for measuring the correctness of PFGs. We present an implementation of SpiderGen, and provide an evaluation of SpiderGen across a set of 66 data points. We find that SpiderGen is able to capture a large portion of required processes for tracking the carbon footprint of various product categories.
Limitations and Open Problems:
Although SpiderGen makes strides towards automating the generation of PFGs, we highlight several limitations of our work that should be addressed to enable workflows like SpiderGen to be adopted in the real world.
Increasing transparency in LCA automation. Recent work indicates that providing transparency in automated LCA methods is a key criteria (ulissi_criteria_2025). Thus, in the design of SpiderGen, we enable transparency by providing sample products, evidence and rationales for processes, and deterministic clustering techniques for forming coarser processes. We hope that this enables human LCA experts to gain a better understanding of how SpiderGen generates PFGs. However, more work must be done to achieve other evaluation goals, such as measuring and indicating the uncertainty of the PFG generation, as well as enabling LCA experts to collaboratively improve the PFG estimates. We hope to explore these possibilities in future work.
Enhancing System Boundaries of PFG Generation. We note that a primary limitation of SpiderGen is derived from the scope, or “system boundary”, of the PFG being in conflict with the system boundary defined appropriately for the product category. In the real world, the system boundary is human-defined. However, for our experiments, we made assumptions about the system boundaries based on ISO standards and encoded them into SpiderGen. We note that this is the root cause for SpiderGen missing processes or hallucinating processes. Although the precision of SpiderGen is relatively high (on average 83%), we note that SpiderGen occasionally produces erroneous processes due to sticking closely to system boundaries that may be inappropriate for a given product. For example, a more common error was to include processes such as “replacement”, “quality control”, or “energy consumption in use” when it was not relevant for a specific product category. We observe in our experiments that SpiderGen misses some the auxiliary processes which are not directly related to the product manufacturing, material composition or usage. We believe that exploring human-AI collaboration methods to make appropriate choices for system boundaries is a promising next step for future work.
Evaluating PFGs in the wild. A primary challenge with evaluating a workflow such as SpiderGen in contexts where there is either no ground-truth available, or where there may be disagreements between experts. For example, even with the ground-truth PCRs that we utilized for our evaluation, there may be experts who disagree on which processes should be included in a PFG for a given product category. An ongoing challenge for future work would be to address enabling expert-evaluation for LCA automation tools, such that expert-consensus is considered.
Downstream Applications: In future work, we hope to further explore the potential of SpiderGen for automating LCA in a variety of contexts, such as for Environmental Product Declarations (EPDs), Material Flow Analysis, which studies the flow of materials through a system, such as an economy, or a manufacturing process.
Acknowledgments
This work is supported by the National Science Foundation Award CNS-2106299, as well as the National Science Foundation Graduate Research Fellowship Program under Grant No(s) (DGE-2140739). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation).